OpenStudy (anonymous):

WHICH NUMBER IS NEXT 45, 90, 135, 180

OpenStudy (anonymous):

225

OpenStudy (anonymous):

you are adding 45 each term

OpenStudy (saifoo.khan):

180 + 45

OpenStudy (anonymous):

225

OpenStudy (ybarrap):

45 = 1*45 90 = 2*45 135 = 3*45 180 = 4*45 Ans 5*45 = 225